OpenStudy (anonymous):

does the mean value theorem apply to |x-1|?

OpenStudy (mathmate):

The mean value theorem applies to a specified closed interval, and it must be differentiable within the (open) interval. This should help you put restriction on whether the MVT applies to |x-1| or not.

OpenStudy (anonymous):

the interval is [0,4]

OpenStudy (anonymous):

... you haven't given us enough information. It must be continuous over \([a,b]\) and differentiable over \((a,b)\), i.e. \(0\notin(a,b)\).

OpenStudy (anonymous):

Ok, you tell us: is \(|x-1|\) continuous on \([a,b]\) and differentiable on \((a,b)\)?

OpenStudy (mathmate):

Hint:  and try to answer questions by oldrin.

OpenStudy (anonymous):

i got to that part but does the vertical asymptote have anything to do with it?

OpenStudy (mathmate):

The main questions are: is f(x) continuous on (0,4), and is f(x) differentiable on [0,4]

OpenStudy (anonymous):

yes to both

OpenStudy (mathmate):

@geerky42 are we able to determine the derivative of f(x)=|x-1| at x=1?

OpenStudy (mathmate):

@abccindy same question for you then: are we able to determine the derivative of f(x)=|x-1| at x=1?

OpenStudy (anonymous):

no because it is 0

OpenStudy (mathmate):

If it is zero, then the answer is yes, because zero is a number.

OpenStudy (anonymous):

it is 1-1/ |1-1|

OpenStudy (mathmate):

Is the derivative really zero?

OpenStudy (anonymous):

\[f'(x)=\frac{ x-1 }{ |x-1| }\]

OpenStudy (mathmate):

For the derivative to exist, the left and right limits of f(x) must be equal.

OpenStudy (mathmate):

...exist and equal.

OpenStudy (anonymous):

The derivative is undefined... c'mon people! There are two different limits of for the slope at x=1 depending on whether you approach from the left or right.

OpenStudy (mathmate):

Would you propose the left limit, i.e. f'(1-), and the right limt f'(1+).

OpenStudy (anonymous):

Yes that is correct.

OpenStudy (mathmate):

So if the derivative is undefined at x=1, what can we conclude if the MVT applies to the interval (0,4)?

OpenStudy (anonymous):

they are all undefined?

OpenStudy (mathmate):

Yes, the derivative at x=1 is undefined because f'(1-)=-1, f'(1+)=+1, so f'(1-)\(\ne\)f'(1+), by definition of the derivative, it is undefined at x=1. This concept is even more important when you work with multi-variable calculus in the near future.